# On the Peril of (Even a Little) Nonstationarity in Satisficing Regret Minimization

### 19 Mar 2026 v1

@@ 6.2 Proof of Theorem 2
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$.
We assume Gaussian rewards of the form $r_{t}(a)=\mu_{t}(a)+\xi_{t}$, where $\xi_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, assume $L$ is even and that $2T/L$ is an integer. Define the block length $m:=2T/L$ and partition the horizon into $L/2$ disjoint blocks $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\},$ $\forall\ell\in[L/2].$ Within each block $\ell$, there is a single change-point $\nu_{\ell}$. Fix integers $n,l\geq 1$ with $nl\leq m-1$, and define the candidate change-point locations (within a block) by $\mathcal{T}:=\{2,2+l,2+2l,\dots,2+(n-1)l\},$ so that the windows $\tau,\tau+l-1]$ are disjoint for distinct $\tau\in\mathcal{T}$. Draw $\nu=(\nu_{1},\dots,\nu_{L/2}),$ where $\nu_{\ell}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}).$ interpreting $\nu_{\ell}$ as the (within-block) change point in block $\ell$, and define the corresponding absolute time $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each time $t\in\mathcal{I}_{\ell}$, set
@@
This construction yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$ be the complete interaction transcript, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{L/2})$ for the collection of all components except the $\ell$-th. Fix $\ell\in[L/2]$ and $t\in[T]$, and condition on $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})=(f,a,x)$. Define
@@
$I(\nu_{\ell};r_{t}\mid F_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $=I\bigl(\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x\bigr)$
@@
If $t\notin\mathcal{I}_{\ell}$, then $r_{t}$ does not depend on $\nu_{\ell}$ given $(\mathcal{F}_{t-1},a_{t},b,u_{-\ell})$, and therefore
$I(\nu_{\ell};r_{t}\mid F_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)=0,\quad\forall t\in[T]\setminus\mathcal{I}_{\ell}.$
@@
$I(\nu_{\ell};r_{t}\mid F_{t-1},a_{t},\nu_{-\ell})$ $\leq 4\Delta^{2}\mathbb{E}\Bigl[\mathbb{P}(\widetilde{\nu}_{\ell}\leq t\mid F_{t-1},a_{t},\nu_{-\ell})\mathds{1}\{a_{t}=2\}+\mathbb{P}(\widetilde{\nu}_{\ell}>t\mid F_{t-1},a_{t},\nu_{-\ell})\mathds{1}\{a_{t}=1\}\Bigr]$
$=4\Delta^{2}\mathbb{E}\Bigl[\mathds{1}\{\widetilde{\nu}_{\ell}\leq t,a_{t}=2\}+\mathds{1}\{\widetilde{\nu}_{\ell}>t,a_{t}=1\}\Bigr].$
@@
$I(u_{\ell};Z\mid B,u_{-\ell})\leq 4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr].$ (17)
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$.
Fix $\ell\in[L/2]$. For each candidate change-point $\tau\in\mathcal{T}$, define
$B_{\ell,\tau}:=\bigl[(\ell-1)m+\tau,\;(\ell-1)m+\tau+l-1\bigr],\qquad N_{\ell,\tau}(Z):=\sum_{t\in B_{\ell,\tau}}\mathds{1}\{a_{t}=1\}.$
Fix an integer $r$ with $1\leq r\leq l$. For each $\tau$ such that $N_{\ell,\tau}(Z)\geq r$, let $\overline{X}_{\ell,\tau}(r)$ denote the average of the first $r$ rewards of arm $1$ observed within $B_{\ell,\tau}$. For each $\tau$ such that $N_{\ell,\tau}(Z)<r$, let $\overline{X}^{\prime}_{\ell,\tau}(l-r)$ denote the average of the first lrl-r rewards of arm $2$ observed within $B_{\ell,\tau}$. Define the estimators
$\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)\geq r\ \text{ and }\ \overline{X}_{\ell,\tau}(r)\geq S\Bigr\}$
$\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)<r\ \text{ and }\ \overline{X}_{\ell,\tau}^{\prime}(l-r)\leq S\Bigr\}$
@@
Fix integers $l,r\in\mathbb{N}$ with $1\leq r\leq l$, and $\ell\in[L/2]$. Then,
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}(N_{\ell,\nu_{\ell}}(Z)<r)+n\exp(-r\Delta^{2}/2),$
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}\big(N_{\ell,u_{\ell}}(Z)\geq r\big)+n\exp(-(l-r)\Delta^{2}/2).$
Construct the final estimator by flipping an independent coin $\Xi\sim\mathrm{Unif}\{0,1\}$. If $\Xi=0$, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$; otherwise, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$. Combining Lemma 2 with the choices $r=\lceil l/2\rceil$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$, we obtain
@@
Combining (17) and (18), we obtain
$4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{1}{4}\log n-\log 2,\quad\forall\ell\in[L/2].$
Consequently,
$\mathbb{E}[R_{T}^{S}(\pi;\nu)]=\Delta\cdot\sum_{\ell\in[L/2]}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\nu_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\nu_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{L}{8\Delta}\Bigl(\frac{1}{4}\log n-\log 2\Bigr).$
Optimize $n$.
Recall the constraints $nl\leq 2T/L-1$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$. Ignoring the ceiling for simplicity, it suffices to choose $n\geq 2$ such that $n\log n\;\leq\;\frac{\Delta^{2}}{6}(\frac{2T}{L}-1).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}/L\geq 13$, we further have $\frac{\Delta^{2}}{6}(\frac{2T}{L}-1)\geq 13/3-1/32>4.$ By Lemma 1, we may choose $n\geq 2$ so that $\mathbb{E}\!\left[R_{T}^{S}(\pi;\nu)\right]\gtrsim\frac{L}{\Delta}\log(\Delta^{2}T/L)$. This completes the proof of Theorem 2.
@@ 6.2.1 Proof of Lemma 2
In this section, we prove the upper bound on $\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ in Lemma 2; the second bound follows by an analogous argument. Let $E:=\{N_{\ell,\nu_{\ell}}(Z)<r\}$. Then
@@
On the event $E^{\mathrm{c}}$, the estimator $\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ can make error only if either
  • the block at the true change-point is not declared, i.e., $\overline{X}_{\ell,\nu_{\ell}}<S$; or
  • some earlier candidate is declared first, i.e., $\overline{X}_{\ell,\tau}\geq S$ for some $\tau<\nu_{\ell}$ with $N_{\ell,\tau}(Z)\geq r$.
For the first event, since $\overline{X}_{\ell,\nu_{\ell}}\sim\mathcal{N}(S+\Delta,1/r)$ on $E^{\mathrm{c}}$, the Gaussian tail bound gives
$\mathbb{P}(\overline{X}_{\ell,\nu_{\ell}}<S)=\mathbb{P}\left(\overline{X}_{\ell,\nu_{\ell}}-(S+\Delta)<-\Delta\right)\leq\exp(-r\Delta^{2}/2).$
Next, fix any $\tau<\nu_{\ell}$. On $E^{\mathrm{c}}\cap\{N_{\ell,\tau}(Z)\geq r\}$, we have $\overline{X}_{\ell,\tau}\sim\mathcal{N}(S-\Delta,1/r)$, and hence
$\mathbb{P}(\overline{X}_{\ell,\tau}\geq S)=\mathbb{P}(\overline{X}_{\ell,\tau}-(S-\Delta)\geq\Delta)\leq\exp(-r\Delta^{2}/2).$
@@
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}})\leq(n-1)\exp(-r\Delta^{2}/2)+\exp(-r\Delta^{2}/2)=n\exp(-r\Delta^{2}/2).$
@@ 6.3 Proof of Theorem 3
Taking a union bound over all $a\in[K]$, epochs $\kappa\leq T$, times $t\leq T$, and windows $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$, and using that $|\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))|\leq T$, we obtain $\mathcal{E}_{\operatorname{win}}^{c}\leq\frac{2}{T}.$ On the event $\mathcal{E}_{\operatorname{win}}$, the windowed confidence bounds behave as follows:
  • Fast rejection. If $\mu_{a,\ell}=S-\Delta$ and $\mathcal{B}_{a}^{(\kappa)}$ contains a suffix window of length $w\geq 2n_{\Delta}$ entirely within segment $\ell$, then $\mathrm{UCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\leq S$. Hence any leader whose mean stays $\leq S-\Delta$ in an epoch is demoted after at most $2n_{\Delta}$ pulls within a single segment.
  • Fast certification. Similarly, any arm whose mean stays $\geq S+\Delta$ in an epoch is certified as satisficing once it accumulates $2n_{\Delta}$ pulls within a single segment.
  • At most one false leader per segment. Within a fixed segment $\ell$, any non-satisficing arm can be certified as leader at most once: after it is demoted its buffer resets, and while its mean remains $\leq S-\Delta$, it cannot satisfy $\mathrm{LCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\geq S$ again before the segment ends.
  • No demotion of the optimal arm. Within segment $\ell$, once an optimal (satisficing) arm is certified as leader, it is never demoted before the segment ends.
On the event $\mathcal{E}_{\operatorname{win}}$, the exploration length on each segment $\ell$ is at most $2Kn_{\Delta_{\ell}^{*}}$: during exploration the algorithm runs a fixed round–robin schedule, and once the optimal arm on segment $\ell$ has been pulled $2n_{\Delta_{\ell}^{*}}$ times, it is certified as satisficing and exploration terminates. Moreover, since every non-satisficing arm $a$ can be certified as leader at most once in segment $\ell$ and, once certified, is demoted after at most $2n_{\Delta_{a,\ell}}$ pulls, the regret incurred on segment $\ell$ is bounded by
$\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log KT\lesssim\sum_{\ell=0}^{L-1}\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T$
up to a universal constant. Therefore, we have

### 25 Apr 2026 v2

@@ 6.2 Proof of Theorem 2
Construct $\{P_{\nu}\}_{\nu\in\mathcal{V}}$. Assume $\xi_{t}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$. For simplicity, suppose $L$ is even and $2T/L\in\mathbb{N}$. Let $m:=\frac{2T}{L}$ and $\mathcal{I}_{\ell}:=\{(\ell-1)m+1,\dots,\ell m\}$ for any $\ell\in[L/2].$ Then the horizon is partitioned into $L/2$ disjoint blocks of length $m$, each containing one change point. Fix integers $n,l\geq 1$ with $nl\leq m-1$, and define $\mathcal{T}:=\{2,\,2+l,\,\dots,\,2+(n-1)l\},$ so that $\tau,\tau+l-1]$ and $\tau^{\prime},\tau^{\prime}+l-1]$ are disjoint for $\tau\neq\tau^{\prime}$. Draw
$\nu=(\nu_{1},\dots,\nu_{L/2}),\qquad\nu_{\ell}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathrm{Unif}(\mathcal{T}),$
and interpret $\nu_{\ell}$ as the within-block change point in block $\ell$. Its absolute location is $\widetilde{\nu}_{\ell}:=(\ell-1)m+\nu_{\ell}\in\mathcal{I}_{\ell}.$ For each block $\ell$ and each $t\in\mathcal{I}_{\ell}$, set
@@
This yields $L$ stationary segments. Let $Z:=(a_{1},r_{1},\dots,a_{T},r_{T})$, and let $P_{\nu}$ denote the law of $Z$ conditional on $\nu$. We write $\nu_{-\ell}:=(\nu_{1},\dots,\nu_{\ell-1},\nu_{\ell+1},\dots,\nu_{L/2})$ for the collection of all components except the $\ell$-th. Fix $\ell\in[L/2]$ and $t\in[T]$, and condition on $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})=(f,a,x)$. Define
@@
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)$ $=I\bigl(\mathds{1}\{\widetilde{\nu}_{\ell}\leq t\};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x\bigr)$
@@
If $t\notin\mathcal{I}_{\ell}$, then $r_{t}$ does not depend on $\nu_{\ell}$ given $(\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$, and therefore
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1}=f,a_{t}=a,\nu_{-\ell}=x)=0,\quad\forall t\in[T]\setminus\mathcal{I}_{\ell}.$
@@
$I(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1},a_{t},\nu_{-\ell})$ $\leq 4\Delta^{2}\mathbb{E}\Bigl[\mathds{1}\{\widetilde{\nu}_{\ell}\leq t,a_{t}=2\}+\mathds{1}\{\widetilde{\nu}_{\ell}>t,a_{t}=1\}\Bigr].$
@@
$\sum_{t=1}^{T}I\bigl(\nu_{\ell};r_{t}\mid\mathcal{F}_{t-1},a_{t},\nu_{-\ell}\bigr)\leq 4\Delta^{2}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\widetilde{\nu}_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\widetilde{\nu}_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr].$ (17)
Construct $\widehat{\nu}_{\ell}(Z,\nu_{-\ell})$ with upper bounded identification error. Fix $\ell\in[L/2]$. For each candidate change-point $\tau\in\mathcal{T}$, define $B_{\ell,\tau}:=\bigl[(\ell-1)m+\tau,\;(\ell-1)m+\tau+l-1\bigr]$ and $N_{\ell,\tau}(Z):=\sum_{t\in B_{\ell,\tau}}\mathds{1}\{a_{t}=1\}.$ Fix an integer $c$ with $1\leq c\leq l$. For each $\tau$ such that $N_{\ell,\tau}(Z)\geq c$, let $\overline{X}_{\ell,\tau}(c)$ denote the average of the first $c$ rewards of arm $1$ observed within $B_{\ell,\tau}$. For each $\tau$ such that $N_{\ell,\tau}(Z)<c$, let $\overline{X}^{\prime}_{\ell,\tau}(l-c)$ denote the average of the first lcl-c rewards of arm $2$ observed within $B_{\ell,\tau}$. Define the estimators
$\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)\geq c\ \text{ and }\ \overline{X}_{\ell,\tau}(c)\geq S\Bigr\}$
$\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$ $:=\min\Bigl\{\tau\in\mathcal{T}:\ N_{\ell,\tau}(Z)<c\ \text{ and }\ \overline{X}_{\ell,\tau}^{\prime}(l-c)\leq S\Bigr\}$
@@
Fix integers $l,c\in\mathbb{N}$ with $1\leq c\leq l$, and $\ell\in[L/2]$. Then,
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}(N_{\ell,\nu_{\ell}}(Z)<c)+n\exp(-c\Delta^{2}/2),$
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ $\leq\mathbb{P}\big(N_{\ell,\nu_{\ell}}(Z)\geq c\big)+n\exp(-(l-c)\Delta^{2}/2).$
Construct the final estimator by flipping an independent coin $\Xi\sim\mathrm{Unif}\{0,1\}$. If $\Xi=0$, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$; otherwise, set $\widehat{\nu}_{\ell}(Z,\nu_{-\ell}):=\widehat{\nu}_{\ell}^{\prime\prime}(Z,\nu_{-\ell})$. Combining Lemma 2 with the choices $c=\lceil l/2\rceil$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$, we obtain
@@
Combining (17) and (18), and using (11), we obtain
$\mathbb{E}[R_{T}^{S}(\pi;\nu)]=\Delta\cdot\sum_{\ell\in[L/2]}\mathbb{E}\Bigl[\sum_{t=(\ell-1)m+1}^{\widetilde{\nu}_{\ell}-1}\mathds{1}\{a_{t}=1\}+\sum_{t=\widetilde{\nu}_{\ell}}^{\ell m}\mathds{1}\{a_{t}=2\}\Bigr]\geq\frac{L}{8\Delta}\Bigl(\frac{1}{4}\log n-\log 2\Bigr).$
Optimize $n$. Recall the constraints $nl\leq 2T/L-1$, $l=\big\lceil 6\log(n)/\Delta^{2}\big\rceil+1$, and $n\geq 2$. Ignoring the ceiling for simplicity, it suffices to choose $n\geq 2$ such that $n\log n\;\leq\;\frac{\Delta^{2}}{6}(\frac{2T}{L}-1).$ Since $\Delta\leq\tfrac{1}{2}$ and $T\Delta^{2}/L\geq 13$, we further have $\frac{\Delta^{2}}{6}(\frac{2T}{L}-1)\geq 13/3-1/32>4.$ By Lemma 1, we may choose $n\geq 2$ so that $\mathbb{E}\!\left[R_{T}^{S}(\pi;\nu)\right]\gtrsim\frac{L}{\Delta}\log(\Delta^{2}T/L)$. This completes the proof of Theorem 2.
@@ 6.2.1 Proof of Lemma 2
In this section, we prove the upper bound on $\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell})$ in Lemma 2; the second bound follows by an analogous argument. Let $E:=\{N_{\ell,\nu_{\ell}}(Z)<c\}$. Then
@@
On the event $E^{\mathrm{c}}$, the estimator $\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})$ can make error only if either the block at the true change point is not declared, i.e., $\overline{X}_{\ell,\nu_{\ell}}<S$; or some earlier candidate is declared first, i.e., $\overline{X}_{\ell,\tau}\geq S$ for some $\tau<\nu_{\ell}$ with $N_{\ell,\tau}(Z)\geq c$.
For the first event, since $\overline{X}_{\ell,\nu_{\ell}}\sim\mathcal{N}(S+\Delta,1/c)$ on $E^{\mathrm{c}}$, the Gaussian tail bound gives $\mathbb{P}(\overline{X}_{\ell,\nu_{\ell}}<S)=\mathbb{P}\big(\overline{X}_{\ell,\nu_{\ell}}-(S+\Delta)<-\Delta\big)\leq\exp(-c\Delta^{2}/2).$
Next, fix any $\tau<\nu_{\ell}$. On $E^{\mathrm{c}}\cap\{N_{\ell,\tau}(Z)\geq c\}$, we have $\overline{X}_{\ell,\tau}\sim\mathcal{N}(S-\Delta,1/c)$, and hence $\mathbb{P}\big(\overline{X}_{\ell,\tau}\geq S\big)=\mathbb{P}\big(\overline{X}_{\ell,\tau}-(S-\Delta)\geq\Delta\big)\leq\exp(-c\Delta^{2}/2).$
@@
$\mathbb{P}(\widehat{\nu}_{\ell}^{\prime}(Z,\nu_{-\ell})\neq\nu_{\ell},E^{\mathrm{c}})\leq(n-1)\exp(-c\Delta^{2}/2)+\exp(-c\Delta^{2}/2)=n\exp(-c\Delta^{2}/2).$
@@ 6.3 Proof of Theorem 3
Taking a union bound over all $a\in[K]$, epochs $\kappa\leq T$, times $t\leq T$, and windows $w\in\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))$, and using that $|\mathrm{WinSet}(\mathcal{B}_{a}^{(\kappa)}(t))|\leq T$, we obtain $\mathbb{P}(\mathcal{E}_{\operatorname{win}}^{c})\leq\frac{2}{T}.$ On the event $\mathcal{E}_{\operatorname{win}}$, the windowed confidence bounds behave as follows:
  • If $\mu_{a,\ell}=S-\Delta$ and $\mathcal{B}_{a}^{(\kappa)}$ contains a suffix window of length $w\geq 2n_{\Delta}$ entirely within segment $\ell$, then $\mathrm{UCB}^{\mathrm{win}}(\mathcal{B}_{a}^{(\kappa)})\leq S$. Hence any leader whose mean stays $\leq S-\Delta$ in an epoch is demoted after at most $2n_{\Delta}$ pulls within a single segment.
  • Similarly, any arm whose mean stays $\geq S+\Delta$ in an epoch is certified as satisficing once it accumulates $2n_{\Delta}$ pulls within a single segment.
  • Within a fixed segment $\ell$, any non-satisficing arm can be certified as leader at most once.
  • Within a fixed segment $\ell$, once an optimal (satisficing) arm is certified as leader, it is never demoted before the segment ends.
On the event $\mathcal{E}_{\operatorname{win}}$, the exploration length on each segment $\ell$ is at most $2Kn_{\Delta_{\ell}^{*}}$: during exploration the algorithm runs a fixed round–robin schedule, and once the optimal arm on segment $\ell$ has been pulled $2n_{\Delta_{\ell}^{*}}$ times, it is certified as satisficing and exploration terminates. Moreover, since every non-satisficing arm $a$ can be certified as leader at most once in segment $\ell$ and, once certified, is demoted after at most $2n_{\Delta_{a,\ell}}$ pulls, the regret incurred on segment $\ell$ is bounded by $\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log KT\lesssim\sum_{a:\mu_{a,\ell}<S}(\frac{1}{\Delta_{a,\ell}}+\frac{\Delta_{a,\ell}}{(\Delta_{\ell}^{*})^{2}}\big)\cdot\log T,$ up to a universal constant. Therefore, we have